OpenStudy (cj49):

There are 30 terms in an arithmetic pregression. The 2nd and 3rd terms are distinct integers. The ratio of the 1st 20 terms and the sum of the 1st 10 terms equals twice the ratio of the 2nd and the 1st terms.Which of the following cn b the sum of all its terms? a.1120 b.1560 c.2020 d.3750

OpenStudy (mokeira):

is something missing from the third sentence

OpenStudy (cj49):

no mate...dis iz d cmplete question

OpenStudy (mokeira):

The ratio of the 1st 20 terms and the sum of the 1st 10 terms equals twice the ratio of the 2nd and the 1st terms the ratio of what of the 1st 20 terms

OpenStudy (cj49):

20/2(2a+19d)/10/2(2a+9d)=2(a+d)/a

OpenStudy (cj49):

the ratio of the "sums"

OpenStudy (mokeira):

ok..give me a min

OpenStudy (anonymous):

if you can form equations from above expression you got, you will get a=9d/8. But any idea what's a or d?

OpenStudy (mokeira):

I think something is missng

OpenStudy (cj49):

a= first no nd d= d difference

OpenStudy (cj49):

so 8a=9d or d=0

OpenStudy (cj49):

no mate dis iz d cmplete question @mokeira

OpenStudy (anonymous):

ok so you can do this thing, go back to your equation substitute a=9d/8 in place of a and try to get the value of "d" out of it.

OpenStudy (cj49):

i stil wnt gt d ans coz equation wud b 30/2(2a+(30-1)d)=x

OpenStudy (anonymous):

from where that x came from? x is what?"

OpenStudy (cj49):

x is the sum of all the 30 terms......i hv to find d sum of all d 30 terms

OpenStudy (anonymous):

yeah that won't help..1 min

OpenStudy (cj49):

20/2(2a+19d)/10/2(2a+9d)=2(a+d)/a=d(8a-9d) d=0 or 8a=9d...dunno hw to proceed after dis

OpenStudy (cj49):

@ganeshie8

OpenStudy (anonymous):

d is looking fine to me..

OpenStudy (cj49):

yups dats d ans

OpenStudy (anonymous):

Till \(8a = 9d\) you got??

OpenStudy (cj49):

yes

OpenStudy (anonymous):

Now, you are given that second and terms are distinct integers.. Note that the words, Distinct Integers, it is an AP so second and third term will be distinct no doubt, but the point to note is, they should be INTEGERS.. and this is the key point here..

OpenStudy (anonymous):

*third is missing in first line.. :)

OpenStudy (cj49):

so d is nt equal to 0

OpenStudy (cj49):

The 2nd and 3rd terms are distinct integers.

OpenStudy (anonymous):

Your progression will be : a, a+d, a+2d, and so on.. Here, a+d is the second term, And a + 2d is the third term.. Right??

OpenStudy (cj49):

yes

OpenStudy (anonymous):

d is the common difference, it must not be 0 for an AP to form..

OpenStudy (anonymous):

I am finding second and third term here..

OpenStudy (anonymous):

First term is : a

OpenStudy (anonymous):

Similarly can you find now third term??

OpenStudy (anonymous):

Wait I have done mistake there.. \[a_2 = a + d = a + \frac{8}{9}a \implies a_2 = \frac{17}{9}a\]

OpenStudy (anonymous):

Now find third term here..

OpenStudy (anonymous):

Can you find third term here in terms of a?

OpenStudy (cj49):

25/9a

OpenStudy (cj49):

it wud b 25/9a ryt

OpenStudy (anonymous):

Yes... Good..

OpenStudy (anonymous):

Now you know what Integers are??

OpenStudy (anonymous):

There @cj49 ??

OpenStudy (cj49):

yesssss d d third term wud b 25/9a??

OpenStudy (anonymous):

Yes I said, you are right..

OpenStudy (anonymous):

For second and third term to be Integer, a must be 9 (minimum) or any multiple of 9..

OpenStudy (cj49):

so dat means d shud b an multiple of 8

OpenStudy (anonymous):

d is already multiple of 8..

OpenStudy (cj49):

30/2(2a+29d)

OpenStudy (anonymous):

See: \[d = \frac{8a}{9}\] Factor 8 is there in numerator which shows d is multiple of 8..

OpenStudy (cj49):

lol yes i missed dat

OpenStudy (anonymous):

So, a = 9 or 18 or 27 likewise multiples of 9.. So, for least value you can take a = 9..

OpenStudy (anonymous):

Just put d = 8 and a = 9 in sum upto 30 terms there..

OpenStudy (cj49):

okkkzzzzzz so D fits d bill...

OpenStudy (anonymous):

Note that if you are taking a = 9, then d = 8 there..

OpenStudy (cj49):

tysm :)....

OpenStudy (anonymous):

Also your question is pretty much clear that it is asking for what CAN BE the sum and not what WILL BE the sum, that means a and d values will change but a will be multiple of 9 only everytime..

OpenStudy (anonymous):

If you have in question given that, what will be the sum, then you will get purely one valid value for a and one for d, but here what can be the sum it is asking, so there can be more values of a and d and hence more values of sum.. Like a = 9 and d = 8 will give d option.. a = 18 and d = 16 will give you some other value.. Likewise.. Hope you are getting what I am trying to say..

OpenStudy (cj49):

its will so i juz hv to substitue d vales nd gt an possible ans nd D is d nli one dat satisfies d cndition

OpenStudy (cj49):

yes i understud dat as d value of a wud change d sum wud also change

OpenStudy (anonymous):

Good question ) @cj49

OpenStudy (anonymous):

@waterineyes what if common difference is mixed number or suppose it is irrational roots? then how we can suppose minimum value?

OpenStudy (anonymous):

We are not concerned with what out d is, you are given that second and third term would be integers and you have those terms as 17a/9 and 25a/9, here to make these terms integers, a must be 9 or multiple of 9, so that denominator 9 will get cancelled and we get a pure integer here.. Now point to note here is, if second term is an integer, suppose let us say second term is 17, can you tell me if d is mixed value, then your third term will definitely going to be mixed which will contradict question's statement.. So, d = integer has to be the case here.. Afterall, if you add an integer to an integer, then only you will get another integer. :)

OpenStudy (anonymous):

ok understood )